# Neural Flow Diffusion Models: Learnable Forward Process for Improved Diffusion Modelling

### v1

@@ A.2 NFDM with restrictions
In Section 4, we discussed training the NFDM with restrictions on the reverse process. This section aims to explore the limitations of such restrictions and to further discuss the training of NFDM with curvature penalties.
@@
When training NFDM with certain restrictions, it is crucial that these restrictions are feasible. For example, imposing overly stringent constraints on the reverse process (such as fixing its drift term $\hat{f}\equiv 0$) will render the forward process incapable of adapting, regardless of its flexibility. Similarly, unattainable penalties, like those on the squared norm of the drift term $\|\hat{f}_{\theta,\varphi}\|_{2}^{2}$, will lead to biased solutions, as it’s impossible to match two processes when $\|\hat{f}_{\theta,\varphi}\|_{2}^{2}\equiv 0$.
Hence, in Section 4.1, we suggest penalizing the curvature of deterministic generative trajectories specifically with penalty ${\mathcal{L}_{\mathrm{crv}}}$ (eq. 26). This is because for non-degenerate data distributions, there exist mappings that transform them into unit Gaussian noise along straight line trajectories.
@@
The second term in eq. 41 is the time derivative of a function. As discussed in section 3.1, the time derivatives can be determined as a Jacobian Vector Product (JVP)with a one-dimensional unit vector. Notably, the first term in eq. 41 is also a JVP. Therefore, we can combine these two operations. For this purpose, we define a $\mathbb{R}^{D+1}$ dimensional adjoint vector $v$:
@@
Consequently, ${\mathcal{L}_{\mathrm{crv}}}$ can be computed as the JVP of $\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)$ with the vector $v$.
Appendix B Connections of NFDM with other approaches
B.1 Special cases of NFDM
@@
Soft Diffusion. Daras et al. (2022) consider a more general case of fixed forward processes, $q(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};C_{t}\mathbf{x},\sigma_{t}I)$, which can be parameterized as $F(\varepsilon,t,\mathbf{x})=C_{t}\mathbf{x}+\sigma_{t}\varepsilon$. Such distributions include, for example, combinations of blurring and the injection of Gaussian noise.
@@
Star-Shaped Diffusion. Okhotin et al. (2023) extended conventional diffusion models to include distributions from the exponential family. Although Star-Shaped Diffusion is a discrete-time approach and does not directly correspond to NFDM, the latter can work with exponential family distributions through reparameterization functions. For instance, for some continuous one-dimensional distribution $q(\mathbf{z}_{t}|\mathbf{x})$, NFDM could use $F(\varepsilon,t,\mathbf{x})=a(b(\varepsilon),t,\mathbf{x})$, where $a$ is the inverse Cumulative Distribution Function (CDF) of $q(\mathbf{z}_{t}|\mathbf{x})$, and $b$ is the CDF of a unit Gaussian.
@@
Variational Diffusion Models. Kingma et al. (2021) proposed forward conditional distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ as $\mathcal{N}(\mathbf{z}_{t};\alpha_{\varphi}(t)\mathbf{x},\sigma_{\varphi}^{2}(t)I)$ with learnable parameters $\varphi$. In the context of NFDM, this distribution can be parameterized by $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{\varphi}(t)\mathbf{x}+\sigma_{\varphi}(t)\varepsilon$.
@@
LSGM. Vahdat et al. (2021) suggest an alternative approach for parameterizing the forward process, proposing diffusion in a latent space of a VAE. Therefore, the forward process is characterized by a distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}E_{\varphi}(\mathbf{x}),\sigma_{t}^{2}I)$, where $E_{\varphi}$ is the encoder of the VAE. To parameterize the same forward process with NFDM, one could use $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{t}E_{\varphi}(\mathbf{x})+\sigma_{t}\varepsilon$. To align the reverse process, the reconstruction distribution should be $p(\mathbf{x}|\mathbf{z}_{0})=\mathcal{N}(\mathbf{x};D_{\varphi}(\mathbf{z}_{0}),\delta^{2}I)$, where $D_{\varphi}$ is VAE’s decoder.
@@
DiffEnc. Nielsen et al. (2023) proposed a more general forward process, $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}f_{\varphi}(\mathbf{x},t),\sigma_{t}^{2}I)$, which, unlike LSGM, transforms $\mathbf{x}$ in a time-dependent manner. This forward process can also be described in terms of NFDM as $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{t}f_{\varphi}(\mathbf{x},t)+\sigma_{t}\varepsilon$.
B.2 Connections of NFDM objective
@@
In this section, we delve into the details of the NFDM’s objective function (eq. 18), with a particular focus on the diffusion loss ${\mathcal{L}_{\mathrm{diff}}}$ (eq. 20).
@@
$=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\frac{1}{2g_{\varphi}^{2}(t)}\left\|\underbrace{\Big{(}f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-f_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}\Big{)}}_{\textrm{Flow Matching term}}+\frac{g_{\varphi}^{2}(t)}{2}\underbrace{\Big{(}\nabla_{\mathbf{z}_{t}}\log q_{\varphi}\big{(}\mathbf{z}_{t}|\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\Big{)}}_{\textrm{Score Matching term}}\right\|_{2}^{2}\right].$ (44)
@@
Crucially, in the extreme scenarios where $g_{\varphi}(t)$ approaches either 00 or $\infty$, the diffusion loss ${\mathcal{L}_{\mathrm{diff}}}$ corresponds to either a reweighted Flow Matching loss (Lipman et al., 2022) or a reweighted Score Matching loss (Vincent, 2011), respectively:
@@
$\lim_{g_{\varphi}(t)\rightarrow\infty}$ ${\mathcal{L}_{\mathrm{diff}}}=\lim_{g_{\varphi}(t)\rightarrow\infty}\mathbb{E}_{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}\left[\frac{g_{\varphi}^{2}(t)}{4}\left\|\nabla_{\mathbf{z}_{t}}\log q_{\varphi}\big{(}\mathbf{z}_{t}|\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\right\|_{2}^{2}\right]$ (46)
B.3 Connections of NFDM with Stochastic Interpolants
In this section, we explore the connections between NFDM and Stochastic Interpolants, as proposed by Albergo et al. (2023). Both NFDM and Stochastic Interpolants introduce a more general family of forward processes through the use of a reparameterization function. Although we acknowledge the relevance of Stochastic Interpolants to our work, there are notable differences between the two approaches.
In their practical applications, Stochastic Interpolants are demonstrated with only simple parameterizations of the forward process. They propose a theoretical framework for learning the forward process that would result in an reverse process characterized by dynamical optimal transport. However, this approach is contingent upon solving a high-dimensional min-max optimization problem, for which they do not provide experimental results. Moreover, their work does not clearly articulate how Stochastic Interpolants might be applied to learning other generative dynamics.
In contrast, NFDM introduces a more generalized method for learning generative dynamics. Moreover, when NFDM is learned with restrictions or penalties (see Section 4), it remains within a min-min optimization paradigm, which can be addressed more efficiently.
Appendix C Implementation details
Our evaluation of NFDM includes tests on synthetic data, CIFAR-10 (Krizhevsky et al., 2009), and two downsampled ImageNet (Deng et al., 2009; Van Den Oord et al., 2016) datasets. To maintain consistency with baselines, we employ horizontal flipping as a data augmentation technique in training models on CIFAR-10 and ImageNet (Song et al., 2020b, 2021). For density estimation of discrete data, uniform dequantization is used (see Section C.1).
We parameterize $\hat{\mathbf{x}}_{\theta}$ (eq. 17) in the reverse process using a 5-layer perceptrons with 512 neurons in each layer for synthetic data and the U-Net architecture from Dhariwal & Nichol (2021) for images. In all experiments, a 3-layer perceptrons with 64 neurons in each layer is employed to parameterize $g_{\varphi}$ (eq. 12), and for $F_{\varphi}$ (eq. 9), we use an identical neural network to that of $\hat{\mathbf{x}}_{\theta}$. The sole difference is that for $F_{\varphi}$, we double the output of the last layer to parameterize $\bar{\mu}_{\varphi}$ (eq. 23) and $\bar{\sigma}_{\varphi}$ (eq. 24) with same model (see Section 3.5).
The models were trained using the Adam optimizer with the following parameters: $\beta_{1}=0.9$, $\beta_{2}=0.999$, a weight decay of $0.0$, and $\epsilon=10^{-8}$. The training process was facilitated by a polynomial decay learning rate schedule, which includes a warm-up phase for a predefined number of training steps. During this phase, the learning rate is linearly increased from $10^{-8}$ to a peak value. After reaching the peak learning rate, it is then linearly decreased to $10^{-8}$ by the final training step. The specific hyperparameters are detailed in Table 3. Training was carried out on Tesla V100 GPUs

### v2

@@ D.1 Additional Discussion of NFDM with Restrictions
In Section 5, we discussed training the NFDM with restrictions on the reverse process. This section aims to further explore the training under restrictions and to examine the limitations of this technique.
Typically, there is an infinite number of forward and reverse processes that correspond to each other. So how to choose between them? The standard NFDM, as described in Section 3, learns just one such pair of processes. However, the flexibility of NFDM opens up the possibility of choosing specific pairs of processes depending on the task.
Suppose we aim to learn a reverse process with some specific properties. To achieve this, we must impose certain restrictions on the reverse process. One approach is to encode these restrictions directly in the parameterization of the reverse process. Alternatively, we can introduce penalties on the reverse process that would penalize deviations from the desired properties. In both cases, we expect the forward process to adapt and align with the reverse process. Consequently, the reverse process will match the forward process and, as a result, the data distribution, while possessing the desired properties.
In Section 5, we discussed the restriction of the reverse process to have straight deterministic generative trajectories. However, we may apply the same strategy to ensure arbitrary properties. We believe that the NFDM framework may help construct various new generative dynamics for a variety of tasks and domains.
@@
Nevertheless, training the NFDM with restrictions has limitations. The primary concern is the feasibility of these restrictions. For instance, imposing overly stringent constraints on the reverse process (such as fixing its drift term $\hat{f}\equiv 0$) will render the forward process incapable of adapting, regardless of its flexibility. Similarly, unattainable penalties, such as those on the squared norm of the drift term $\|\hat{f}_{\theta,\varphi}\|_{2}^{2}$, will lead to biased solutions, as it is impossible to match two processes when $\|\hat{f}_{\theta,\varphi}\|_{2}^{2}\equiv 0$.
Additionally, even when a restriction is feasible, if the model parameterization is too simplistic, it may not be able to match processes and satisfy the restriction, leading to bias. However, in our experiments, we find that even the Gaussian parameterization of the forward process (see Section B.5) is flexible enough in practice to successfully learn various restrictions, such as straight generative trajectories.
D.2 Details of NFDM with Curvature Penalty
In Section 5, we suggest penalizing the curvature of deterministic generative trajectories specifically with the ${\mathcal{L}_{\mathrm{crv}}}$ (eq. 16) penalty.
@@ D.2 Details of NFDM with Curvature Penalty
The second term in eq. 65 is the time derivative of a function. As discussed in section B.2, the time derivatives can be determined as a Jacobian Vector Product (JVP)with a one-dimensional unit vector. Notably, the first term in eq. 65 is also a JVP. Therefore, we can combine these two operations. For this purpose, we define a $\mathbb{R}^{D+1}$ dimensional adjoint vector $v$:
@@
Consequently, ${\mathcal{L}_{\mathrm{crv}}}$ can be efficiently computed as the JVP of $\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)$ with the vector $v$.
Additionally, we would like to note a couple of aspects of learning the NFDM with a curvature penalty. First, unlike the regular NFDM, the curvature penalty necessitates that $F_{\varphi}$ be twice differentiable with respect to time. Second, although we parameterize $\hat{f}_{\theta,\varphi}$ (eq. 11) through $\tilde{f}^{B}_{\varphi}$ (eq. 10), ${\mathcal{L}_{\mathrm{crv}}}$ influences the forward process but does not penalize the curvature of the conditional forward trajectories (see illustrations in Section G.1).
Appendix E Connections of NFDM with Other Approaches
In this section we continue discussing connections between NFDM and related works.
E.1 Special Cases of NFDM
@@
Soft Diffusion. [9] consider the case of a fixed forward processes $q(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};C_{t}\mathbf{x},\sigma_{t}I)$, which can be reparameterized as $F(\varepsilon,t,\mathbf{x})=C_{t}\mathbf{x}+\sigma_{t}\varepsilon$. Such distributions include, for example, combinations of blurring and the injection of Gaussian noise.
@@
Star-Shaped Diffusion. [40] extended conventional diffusion models to include distributions from the exponential family. Although Star-Shaped Diffusion is a discrete-time approach and does not directly correspond to NFDM, the latter can work with exponential family distributions through reparameterization functions. For instance, for some continuous one-dimensional distribution $q(\mathbf{z}_{t}|\mathbf{x})$, NFDM could use $F(\varepsilon,t,\mathbf{x})=a(b(\varepsilon),t,\mathbf{x})$, where $a$ is the inverse Cumulative Distribution Function (CDF) of $q(\mathbf{z}_{t}|\mathbf{x})$, and $b$ is the CDF of a unit Gaussian.
@@
Variational Diffusion Models. [27] proposed forward conditional distributions $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ as $\mathcal{N}(\mathbf{z}_{t};\alpha_{\varphi}(t)\mathbf{x},\sigma_{\varphi}^{2}(t)I)$ with learnable parameters $\varphi$. In the context of NFDM, this can be realized by $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{\varphi}(t)\mathbf{x}+\sigma_{\varphi}(t)\varepsilon$.
@@
LSGM. [65] suggest an alternative approach for parameterizing the forward process, proposing diffusion in a latent space of a VAE. Therefore, the forward process is characterized by a distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}E_{\varphi}(\mathbf{x}),\sigma_{t}^{2}I)$, where $E_{\varphi}$ is the encoder of the VAE. To parameterize the same forward process with NFDM, one could use $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{t}E_{\varphi}(\mathbf{x})+\sigma_{t}\varepsilon$. To align the reverse process, the reconstruction distribution should be $p(\mathbf{x}|\mathbf{z}_{0})=\mathcal{N}(\mathbf{x};D_{\varphi}(\mathbf{z}_{0}),\delta^{2}I)$, where $D_{\varphi}$ is VAE’s decoder.
@@
NDM and DiffEnc. [4, 39] proposed a more general forward process, $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}f_{\varphi}(\mathbf{x},t),\sigma_{t}^{2}I)$, which, unlike LSGM, transforms $\mathbf{x}$ in a time-dependent manner. This forward process can also be described in terms of NFDM as the special case $F_{\varphi}(\varepsilon,t,\mathbf{x})=\alpha_{t}f_{\varphi}(\mathbf{x},t)+\sigma_{t}\varepsilon$.
Curvature minimization. [31] suggested learning the distribution $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})$ of the forward process at time step $t=1$ as a conditional Gaussian. They parameterize the forward process via linear interpolation between $\mathbf{x}$ and $\mathbf{z}_{1}$. Within the NFDM framework, we can express this as $F_{\varphi}(\varepsilon,t,\mathbf{x})=(1-t)\mathbf{x}+t\big{(}\mu_{\varphi}(\mathbf{x})+\sigma_{\varphi}(\mathbf{x})\varepsilon\big{)}$. Subsequently, the authors train their model by optimizing the Flow Matching objective in conjunction with minimizing the Kullback–Leibler (KL) divergence between $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})$ and $p(\mathbf{z})$. This objective corresponds to minimization of diffusion term ${\mathcal{L}_{\mathrm{diff}}}$ and prior term ${\mathcal{L}_{\mathrm{prior}}}$ of NFDM objective (see Section A.1) for the forward process proposed by [31].
E.2 Compatible and Orthogonal Approaches
Despite the simulation-free nature of the training procedure, diffusion models still necessitate full reverse process simulations for sample generation, leading to slow and computationally expensive inference. To address this, in an orthogonal line of works alternative sampling methods have been studied, such as deterministic sampling in discrete [56] and continuous time [60] and novel numerical solvers [61, 34, 37, 52]. Additionally, distillation techniques have been applied to both discrete [49] and continuous time models [57, 73, 71] to enhance sampling speed, albeit at the cost of training additional models. Most of these approaches are compatible with NFDM and may be adapted for further gains.
The significance of deterministic trajectories in efficient sampling was highlighted by [23]. At the same time, building on diffusion model concepts, [32, 36, 2] introduced simulation-free methods for learning deterministic generative dynamics. Based on these ideas, [36] and [35] proposed distillation procedures to straighten deterministic generative trajectories. Since these methods are based on distillation, we consider them orthogonal to NFDM.
E.3 Other Related Works
Schrödinger Bridge models. Recent studies [72, 10, 53] have explored generative models based on Schrödinger Bridge theory and finite-time diffusion constructions. Unlike NFDM, these models deviate from the simulation-free paradigm by necessitating the full simulation of stochastic processes for training, which significantly increases their computational cost. Moreover, these studies typically aim to address the optimal stochastic control problem, which is outside the scope of our work.
Nevertheless, it is important to highlight that within this research domain, some authors have proposed advanced techniques for parameterizing learnable conditional processes. For instance, in [33], the authors introduced the use of splines to parameterize the conditional process, thereby constructing a more flexible and efficient model. This parameterization technique can be seen as a specific instantiation of the parameterization procedure proposed by the NFDM framework.
Minibatch optimal transport. [43] and [63] proposed to construct the forward process with optimal data-noise couplings to learn straighter generative trajectories. While this method is theoretically justified, it relies on estimating an optimal coupling over minibatches of the entire dataset, which, for large datasets, may become uninformative as to the true optimal transport coupling. In contrast, NFDM-OT directly penalizes curvature of the generative trajectories.
Stochastic Interpolants. In this section, we explore the connections between NFDM and Stochastic Interpolants, as proposed by [1]. Both NFDM and Stochastic Interpolants introduce a more general family of forward processes through the use of a reparameterization function. Although we acknowledge the relevance of Stochastic Interpolants to our work, there are notable differences between the two approaches.
In their practical applications, Stochastic Interpolants are demonstrated with a simple parameterizations of the forward process. They propose a theoretical framework for learning the forward process that would result in an reverse process characterized by dynamical optimal transport. However, this approach is contingent upon solving a high-dimensional min-max optimization problem, for which they do not provide experimental results. Moreover, their work does not clearly articulate how Stochastic Interpolants might be applied to learning other generative dynamics.
In contrast, NFDM introduces a more generic method for learning generative dynamics. Moreover, when NFDM is learned with restrictions or penalties (see Section 5), it remains within a min-min optimization paradigm, which can be addressed more efficiently.
E.4 Connections of NFDM Objective
@@
In this section, we delve into the details of the NFDM’s objective function (eq. 12). Specifically we will consider the diffusion term of the objective ${\mathcal{L}_{\mathrm{diff}}}$ (see Section A.1).
@@
$=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}$ $\left[\frac{1}{2g_{\varphi}^{2}(t)}\left\|\underbrace{\Big{(}f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-f_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}\Big{)}}_{\textrm{Flow Matching term}}+\right.\right.$ (68)
$\quad\left.\left.\frac{g_{\varphi}^{2}(t)}{2}\underbrace{\Big{(}\nabla_{\mathbf{z}_{t}}\log q_{\varphi}\big{(}\mathbf{z}_{t}|\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\Big{)}}_{\textrm{Score Matching term}}\right\|_{2}^{2}\right].$
@@
Crucially, in the extreme scenarios where $g_{\varphi}(t)$ approaches either 00 or $\infty$, the diffusion loss ${\mathcal{L}_{\mathrm{diff}}}$ corresponds to either a reweighted Flow Matching loss [32] or a reweighted Score Matching loss [67], respectively:
@@
$\lim_{g_{\varphi}(t)\rightarrow\infty}$ ${\mathcal{L}_{\mathrm{diff}}}=\lim_{g_{\varphi}(t)\rightarrow\infty}\mathbb{E}_{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}\left[\frac{g_{\varphi}^{2}(t)}{8}\left\|\nabla_{\mathbf{z}_{t}}\log q_{\varphi}\big{(}\mathbf{z}_{t}|\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)}-\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})\right\|_{2}^{2}\right]$ (70)
This connection highlights the importance of optimizing the full objective in eq. 67, which contains both Flow Matching and Score Matching components. Simplifying the objective to include only one component renders it intractable from a variational perspective. Specifically, when $g_{\phi}(t)$ approaches either 00 or $\infty$, the objective, a variational bound on likelihood, becomes infinitely large.
There are two primary reasons why many popular diffusion models are trained using just one of these components.
Firstly, many diffusion models employ a simple linear parameterization of the forward process. For these models, both the conditional vector field $f_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})$ and the conditional score function $\nabla_{\mathbf{z}_{t}}\log q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ are merely linear combinations of $\mathbf{z}_{t}$ and $\mathbf{x}$. As [26] demonstrated, for such models, Flow Matching and Score Matching can be viewed as reweighted ELBO. However, for more general non-linear forward process in NFDM, optimizing Flow Matching and Score Matching objectives is not equivalent.
Secondly, in other models, the Flow Matching and Score Matching objectives are derived under the assumption that the forward process is fixed. Without this assumption, we cannot derive either objective in the general case. In NFDM, we learn the forward process end-to-end with the reverse process. Thus, reducing the NFDM’s objective could lead to the model’s collapse, as it is not the correct objective to optimize.
We would also like to emphasize that the NFDM framework, including its objective in eq. 67, represents a generalization of conventional diffusion models, rather than an alternative approach to training. Consequently, if we take an existing forward process from conventional diffusion models and incorporate it into the objective $\mathcal{L}$, we will not create a new model or optimization procedure. Instead, this will align with approaches that focus on likelihood-based training of conventional diffusion models, as explored in [58, 74].
NFDM facilitates the convenient pre-specification of fixed forward processes. Hence, it might be utilized to define a fixed forward process and subsequently train it with the Flow Matching objective. However, this is a topic for future research and is beyond the scope of this work.
Appendix F Implementation Details